OpenStudy (anonymous):

Can anyone help me understand how to write the standard form equation of an ellipse from the vertices and co-vertices? Greatly appreciated.

OpenStudy (anonymous):

Can you explain a little more? Do you mean that the vertices of the major and minor axes are given and you want to write in the following form: $$\frac{x^2}{a^2} + \frac{y^2}{b^2} = 1$$ ??

OpenStudy (anonymous):

As above, with vertex a and covertex b (a >b)

OpenStudy (anonymous):

I'm given the vertex and co-vertex, I need to write the equation, I do not understand how to write the equation given only these two pieces of information.

OpenStudy (anonymous):

It's right there, we just posted it above....

OpenStudy (anonymous):

 Your question is still unclear, leading to most of the confusion I guess. If your question is that you have the coordinates (x_1, y_1) and (x_2, y_2) in the above figure, then just replace x_1 for a and y_2 for b in the above equation.